OpenStudy (anonymous):

y + 2x = −1 3y − x = 4

hartnn (hartnn):

Multiply 2 to 2nd equation and then add both the equations what do u get ?

OpenStudy (anonymous):

To solve this, solve on equation in terms of a single variable, then plug it into the other equation. From there, find the numerical value for the other variable and use that to solve again for the original equation. In other words, solve for "x" in the first equation, and since it's equal to "x", you can put it into the "x" in the second equation. If you can't figure it out from there, I can walk you though the rest.

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

So, for the first equation, let's solve for x: y + 2x = -1 Subtract y from both sides: 2x = -1 - y And divide by two: \[x = \frac{ (-1) -y }{ 2 }\] The stuff on the right side is equal to "x", so we can plug it into the "x" that's in the second equation: 3y − x = 4 \[3y − \frac{ (-1) -y }{ 2 }\ = 4 \] Now you can solve for "y" like your normally would to get an explicit numerical value for it. Can you figure out the rest?

OpenStudy (anonymous):

so then what do you do?

OpenStudy (anonymous):

Well now you have an exact value for "y". Where can you use this to find the exact value of "x"?

OpenStudy (anonymous):

so it'll me (-1,1)?